# Discrete dark matter with light Dirac neutrinos

### v2

@@ IV Dark matter
In this model, gauge singlet Dirac fermions $N_{2,3}$, gauge singlet scalars $\chi_{2,3}$ and $SU(2)$ doublet scalars $\eta_{2,3}$, remain odd under the residual $\mathcal{Z}_{2}$ symmetry such that the lightest among them becomes a viable dark matter candidate. When considering the fermionic dark matter scenario, $i.e.$, $M_{N}<M_{\eta}<M_{\chi}$, achieving the correct relic density becomes challenging because the mass $M_{N}$ is constrained to the multi-TeV range by the neutrino mass criteria. Fermion DM in scotogenic model was studied in several earlier works, see Ahriche:2017iar ; Borah:2018smz ; Mahanta:2019gfe for example. As can be seen from these earlier works, fermion singlet DM relic relies primarily on Yukawa portal annihilations and coannihilations. Due to the involvement of the same Yukawa couplings in both type-I and scotogenic contributions to neutrino mass, we can not have Yukawa coupling in $\geq\mathcal{O}(1)$ regime as discussed in the previous section. After including coannihilation among fermion DM $N$ and $\mathcal{Z}_{2}$-odd scalars, correct relic abundance is found only for Case-A with $M_{N}\in[3000,5000]$ GeV, considering the parameter space in agreement with neutrino data. However, for Case-B ($M_{N}\in[6000,10000]$ GeV) and Case-C ($M_{N}\in[11400,14000]$ GeV), where $M_{N}$ is pushed to even larger values, it is not possible to achieve the correct relic density for the fermionic dark matter via thermal freeze-out. This is because the large $M_{N}$ value and insufficient (co)annihilation rates always lead to an overabundant relic density. Thermal relic of $N$ in Case-A is obtained due to smaller $M_{N}$ and large co-annihilation rate with $SU(2)$ doublet scalars via large Yukawa portal ($y_{i}\gg y^{\prime}_{i}$) interactions. Since doublet scalars have more efficient annihilation rates, it is easier to obtain correct relic of $N$ in Case-A compared to Case-B where $N$ has sizeable Yukawa with scalars which are predominantly singlets. While Case-C offers the possibility of non-thermal or freeze-in fermion singlet DM, we find it to be overproduced as well leaving only Case-A as viable for fermion DM scenario.
@@
Figure 8: Left panel: variation of scalar DM relic abundance with its mass. The blue solid line indicates the observed DM relic. Right panel: spin-independent DM-nucleon scattering as a function of the DM mass. The red colored points satisfy DM relic. The cyan shaded region is ruled out by direct detection constraints.
@@
In the case of scalar DM, the correct relic density can be obtained in all the three cases discussed in the previous section due to the gauge and scalar portal annihilation and co-annihilation of the scalar DM while successfully explaining the neutrino data. Due to the possibility of additional portals namely gauge and scalar portals for annihilation, scalar DM relic can be obtained without relying on the Yukawa portal, unlike fermion singlet DM. With $M_{N}$ varied in the required range so as to satisfy the neutrino oscillation data, we vary the masses of $\mathcal{Z}_{2}$-odd scalars $\zeta_{i}$ (see Appendix B for the details of physical scalar masses and eigenstates) assuming two hierarchies $M_{\zeta_{i}}<M_{N}$ or $M_{\zeta_{1}}<M_{N}<M_{\zeta_{j}},(j\neq 1)$. We consider $\zeta_{1}$, the lightest $\mathcal{Z}_{2}$-odd scalar to be the DM. As the mass difference between $\zeta_{i}$’s are constrained in a specific range from the neutrino mass criteria, when the mass hierarchy is $M_{\zeta_{i}}<M_{N}$, it is possible to get DM masses in the sub-TeV range. However when we assume the hierarchy $M_{\zeta_{1}}<M_{N}<M_{\zeta_{j}}(j\neq 1)$, the DM mass always lies in the multi-TeV scale.
@@
In the left panel of Fig. 8, we show the variation of scalar DM relic with its mass. As shown in several earlier works on scalar DM, the (co)annihilations mediated by electroweak gauge bosons and the SM Higgs are responsible for the dip for intermediate mass, leaving two distinct mass ranges satisfying correct relic. The right panel of the same figure shows the spin-independent DM-nucleon scattering cross-section (mediated via Higgs portal) as a function of scalar DM mass for the points which satisfy relic abundance criteria. We used the package MicrOMEGAs Alguero:2023zol to calculate the relic density and DM-nucleon scattering cross-section after generating the model files using LanHEP Semenov:2014rea . The cyan shaded region is ruled out by the most stringent direct search constraints from LZ-2022 LUX-ZEPLIN:2022qhg for heavier DM and CRESST-III CRESST:2019jnq and DarkSide-50 DarkSide:2018bpj for lighter DM. Clearly we can see that the viable DM mass range that satisfies correct relic density and direct detection rate ranges from around $500$ GeV to several TeV scale. Interestingly, future direct detection experiment Darwin DARWIN:2016hyl can probe a part of this high mass region up to $M_{\rm DM}\sim 3$ TeV.
@@ V Observable $\Delta N_{\rm eff}$
at $2\sigma$ or $95\%$ C.L. including baryon acoustic oscillation (BAO) data. At $1\sigma$ C.L. it becomes more stringent to ${\rm N}_{\rm eff}=2.99\pm 0.17$. Similar bound also exists from big bang nucleosynthesis (BBN), $3.42.3<{\rm N}_{\rm eff}<$ 3.4 at $95\%$ C.L. Cyburt:2015mya . Recent observations by the DESI collaboration report $\Delta N_{\rm eff}<0.4$ at $2\sigma$ DESI:2024mwx , which is weaker than the PLANCK bound given in Eq. (14). All these bounds are consistent with SM predictions ${\rm N^{SM}_{eff}}=3.045$ Mangano:2005cc ; Grohs:2015tfy ; deSalas:2016ztq . Future CMB experiment CMB-S4 is expected reach a much better sensitivity of $\Delta{\rm N}_{\rm eff}={\rm N}_{\rm eff}-{\rm N}^{\rm SM}_{\rm eff}=0.06$ Abazajian:2019eic , taking it closer to the SM prediction. Another future CMB experiment, specially, CMB-HD CMB-HD:2022bsz can probe $\Delta N_{\rm eff}$ up to $0.014$ at $1\sigma$. Dirac nature of light neutrinos introduces additional relativistic degrees of freedom in the form of right chiral neutrinos which can contribute to $\Delta{\rm N}_{\rm eff}$ either thermally or non-thermally. Light Dirac neutrino models often lead to enhanced $\Delta{\rm N}_{\rm eff}$, some recent works on which can be found in Abazajian:2019oqj ; FileviezPerez:2019cyn ; Nanda:2019nqy ; Han:2020oet ; Luo:2020sho ; Borah:2020boy ; Adshead:2020ekg ; Luo:2020fdt ; Mahanta:2021plx ; Du:2021idh ; Biswas:2021kio ; Borah:2022obi ; Li:2022yna ; Das:2023yhv . Depending upon the strength of Yukawa couplings associated with $\nu_{R}$, we can have either thermal or non-thermal contribution to $\Delta{\rm N}_{\rm eff}$ as we discuss in the following subsections.
V.1 Thermal $\Delta N_{\rm eff}$
@@
We have three generations of right handed neutrinos ($\nu_{R_{i}}$) in our model, and their abundance in the Universe will be decided by the interaction strength and the masses of the particle with which they interact. The last three interaction terms of the Lagrangian (Eq. (II)) are the only three interactions that $\nu_{R}$ have with $N$ and $\chi$. The relative strength between the two types of Yukawa couplings $y_{i}$ and $y_{i}^{\prime}$ leads to three distinctive classes (Case-A, Case-B, Case-C) as we discussed before, in the context of neutrino mass and DM. With sufficiently large Yukawa coupling ($y_{i}^{\prime}$) as in Case-B, $\nu_{R}$ may enter the thermal bath with the SM degrees of freedom. As the Universe expands and the temperature falls, the scattering processes that kept $\nu_{R}$ in the thermal bath eventually go out of equilibrium, leading to the decoupling of $\nu_{R}$. For the combined scattering processes of $N_{L}\bar{\nu_{R}}\leftrightarrow N_{L}\bar{\nu_{R}}$ and $\chi\bar{\nu_{R}}\leftrightarrow\chi\bar{\nu_{R}}$, we check the decoupling or thermalization profile with two sets of benchmark parameters (masses and couplings) as shown in Fig. 9. As $N$ and $\chi$ masses lies at $\mathcal{O}$(TeV) range, it leads to the decoupling of $\nu_{R}$ from the thermal bath very early. Such early decoupling, above the electroweak scale, leads to a fixed value $\Delta N_{\rm eff}=0.14$ for three generations of $\nu_{R}$. Clearly, this remains within reach of several future CMB experiments mentioned earlier.
V.2 Non-thermal $\Delta N_{\rm eff}$
It is also possible to have observable $\Delta N_{\rm eff}$ when $\nu_{R}$ does not enter the thermal bath. In such a case, $\nu_{R}$ can be produced non-thermally from the decay of the scalar $\chi$ as $\chi\rightarrow N\nu_{R_{i}}$ or from decay of $N$ as $N\rightarrow\chi\nu_{R_{i}}$ depending upon our choice of fermion or scalar DM. Similar contributions can also arise from scatterings. With small Yukawa couplings for the $\bar{N_{L}}\chi\nu_{R}$ interaction (as in Case-A and Case-C), $\nu_{R}$ can not enter the thermal bath via the $N_{L}\bar{\nu_{R}}\leftrightarrow N_{L}\bar{\nu_{R}}$ (or $\chi\bar{\nu_{R}}\leftrightarrow\chi\bar{\nu_{R}}$) scattering process. In such a situation, where the scalar $\chi$ is in the thermal bath, $\nu_{R}$ are produced via the decay $\chi\rightarrow N\nu_{R}$ (assuming fermion DM scenario).
@@ V.2 Non-thermal $\Delta N_{\rm eff}$
where $\Gamma_{\chi}$ denotes the partial decay width for the process $\chi\rightarrow N\nu_{R}$. The $\nu_{R}$ produced from $\chi$ decay will contribute to this additional radiation density. Therefore, with $\rho_{R}=\rho_{\rm SM}+\rho_{\nu_{R}}$, $\Delta N_{\rm eff}$ for one species of $\nu_{R}$ (indicated by flavour index $\alpha$) calculated at the CMB epoch $(T_{\rm CMB})$ is
$\Delta N^{\alpha}_{\rm eff}$ $=$ $N^{\alpha}_{\rm eff}-N_{\rm eff}^{\rm SM}$ (17)
@@
where, the extrapolation from the BBN epoch $(T_{\rm BBN})$ to CMB epoch $(T_{\rm CMB})$ is justified by the fact that both $\rho_{\nu_{L}},\rho_{\nu_{R}}$ redshift similarly for $T<T_{\rm BBN}$. One can simply calculate the energy density ratio at BBN epoch $\sim\mathcal{O}({10})$ MeV as $\nu_{R}$ production stops much earlier. The numerical factor 2 in Eq. (17) indicate the degrees of freedom of each $\nu_{R}$. The net contribution to $\Delta N_{\rm eff}$ can be obtained by summing over flavour index $\alpha$ as $\Delta N_{\rm eff}=\sum_{\alpha}\Delta N^{\alpha}_{\rm eff}$. Ignoring the temperature dependence of the relativistic degrees of freedom, one can solve Eq. (15) analytically to find $\Delta N_{\rm eff}$ Biswas:2022vkq .
@@
The decoupling profile for $\nu_{R}$ for two different benchmark choices of Yukawa coupling $y_{i}^{\prime}$ is shown in Fig. 9. For scalar ($\chi$) and fermion ($N$) masses in the TeV ballpark, we find that for $y_{i}^{\prime}\leq\mathcal{O}(10^{-3})$, the interaction $N_{L}\bar{\nu_{R}}\leftrightarrow N_{L}\bar{\nu_{R}}$ can not attain thermal equilibrium in the early Universe. Therefore, Case-A with $y_{i}^{\prime}\sim\mathcal{O}(10^{-11})$, $\nu_{R}$ will never enter the thermal bath, hence, thermal $\Delta N_{\rm eff}$ is not possible. On the other hand, $\nu_{R}$ produced non-thermally from the decay of either $\chi$ or $N$ gives rise to negligible contribution to $\Delta N_{\rm eff}$. As mentioned before, the Case-B, with $0.01\leq y_{1}^{\prime}\leq 0.1$, $0.03\leq y_{2}^{\prime}\leq 0.1$, $0.03\leq y_{3}^{\prime}\leq 0.1$, we have thermal contribution to $\Delta N_{\rm eff}$. Due to the decoupling of $\nu_{R}$ much above the electroweak scale, we have a fixed contribution $\Delta N_{\rm eff}=0.14$ for Case-B. Finally, in Case-C, Yukawa couplings are of the size $\mathcal{O}(10^{-6})$, and the corresponding interactions do not allow $\nu_{R}$ to enter the thermal bath. Considering the decay of $\chi\rightarrow\nu_{R}N$, we have shown the non-thermal contributions to $\Delta N_{\rm eff}$ in Fig. 10. Clearly, some part of the parameter space is already ruled out by PLANCK 2018 and DESI 2024 limits while most of the remaining parameter space remains within reach of future CMB experiments like SPT-3G, CMB-S4 and CMB-HD. As there are multiple $\mathcal{Z}_{2}$-odd scalars, it is possible to have $\chi_{i}\rightarrow\nu_{R}N$ type decays from one of the heavier scalars irrespective of fermion or scalar DM. While contribution to $\Delta N_{\rm eff}$ from scalar decay can be calculated simply by considering the scalar $\chi_{i}$ to be in equilibrium, contributions from the decay $N\rightarrow\nu_{R}\chi_{j}$ will be more involved as the number density of non-thermal $N$ needs to be found by solving the respective Boltzmann equation Biswas:2022vkq .
@@
To summarise, observable $\Delta N_{\rm eff}$ is possible in Case-B and Case-C. However, fermion DM is not possible in these two cases irrespective of freeze-out or freeze-in mechanism. On the other hand, fermion DM produced via freeze-out mechanism is viable in Case-A even though the $\Delta N_{\rm eff}$ remains out of reach from future CMB experiments. This conflict between observable $\Delta N_{\rm eff}$ and DM relic can be avoided if we consider scalar DM. Scalar DM allows light scalars in sub-TeV ballpark with observable direct detection rates while keeping collider prospects more promising compared to fermion DM scenario where all $\mathcal{Z}_{2}$-odd particles remain in multi-TeV ballpark.

### v3

@@ IV Dark matter
In this model, gauge singlet Dirac fermions $N_{2,3}$, gauge singlet scalars $\chi_{2,3}$ and $SU(2)$ doublet scalars $\eta_{2,3}$, remain odd under the residual $\mathcal{Z}_{2}$ symmetry such that the lightest among them becomes a viable dark matter candidate. When considering the fermionic dark matter scenario, $i.e.$, $M_{N}<M_{\eta}<M_{\chi}$, achieving the correct relic density becomes challenging because the mass $M_{N}$ is constrained to the multi-TeV range by the neutrino mass criteria. Fermion DM in the scotogenic model was studied in several earlier works; see Ahriche:2017iar ; Borah:2018smz ; Mahanta:2019gfe for example. As can be seen from these earlier works, the fermion singlet DM relic relies primarily on Yukawa portal annihilations and coannihilations. Due to the involvement of the same Yukawa couplings in both type-I and scotogenic contributions to neutrino mass, we can not have Yukawa coupling in $\geq\mathcal{O}(1)$ regime as discussed in the previous section. After including coannihilation among fermion DM $N$ and $\mathcal{Z}_{2}$-odd scalars, correct relic abundance can be found only for Case-A with $3000\leq(M_{N}/\text{GeV})\leq 5000$, considering the parameter space in agreement with neutrino data. However, for Case-B with $6000\leq(M_{N}/\text{GeV})\leq 10000$ and Case-C with $11400\leq(M_{N}/\text{GeV})\leq 14000$, where $M_{N}$ is pushed to even larger values, it is not possible to achieve the correct relic density for the fermionic dark matter via thermal freeze-out. This is because the large $M_{N}$ value and insufficient (co)annihilation rates always lead to an overabundant relic density. The thermal relic of $N$ in Case-A can be obtained due to smaller $M_{N}$ and large co-annihilation rate with $SU(2)$ doublet scalars via large Yukawa portal ($y_{i}\gg y^{\prime}_{i}$) interactions. Since doublet scalars have more efficient annihilation rates, it is easier to obtain the correct relic of $N$ in Case-A compared to Case-B, where $N$ has sizeable Yukawa with scalars which are predominantly singlets. While Case-C offers the possibility of non-thermal or freeze-in fermion singlet DM, we find it to be overproduced as well leaving only Case-A as viable for fermion DM scenario.
@@
Figure 5: Left panel: variation of scalar DM relic abundance with its mass. The blue solid line indicates the observed DM relic. Right panel: spin-independent DM-nucleon scattering as a function of the DM mass. The red coloured points satisfy DM relic. The cyan-shaded region is ruled out by direct detection constraints.
@@
In the case of scalar DM, the correct relic density can be obtained in all three cases discussed in the previous section due to the gauge and scalar portal annihilation and co-annihilation of the scalar DM while successfully explaining the neutrino data. Due to the possibility of additional portals namely gauge and scalar portals for annihilation, scalar DM relic can be obtained without relying on the Yukawa portal, unlike fermion singlet DM. With $M_{N}$ varied in the required range so as to satisfy the neutrino oscillation data, we vary the masses of $\mathcal{Z}_{2}$-odd scalars $\zeta_{i}$ (see Appendix B for the details of physical scalar masses and eigenstates) assuming two hierarchies $M_{\zeta_{i}}<M_{N}$ or $M_{\zeta_{1}}<M_{N}<M_{\zeta_{j}},(j\neq 1)$. We consider $\zeta_{1}$, the lightest $\mathcal{Z}_{2}$-odd scalar to be the DM. As the mass difference between $\zeta_{i}$’s are constrained in a specific range from the neutrino mass criteria when the mass hierarchy is $M_{\zeta_{i}}<M_{N}$, it is possible to get DM masses in the sub-TeV range. However when we assume the hierarchy $M_{\zeta_{1}}<M_{N}<M_{\zeta_{j}}(j\neq 1)$, the DM mass always lies in the multi-TeV scale.
@@
In the left panel of Fig. 5, we show the variation of scalar DM relic with its mass. As shown in several earlier works on scalar DM, the (co)annihilations mediated by electroweak gauge bosons and the SM Higgs are responsible for the dip for intermediate mass, leaving two distinct mass ranges satisfying the correct relic. The right panel of the same figure shows the spin-independent DM-nucleon scattering cross-section (mediated via the Higgs portal) as a function of scalar DM mass for the points which satisfy relic abundance criteria. We used the package MicrOMEGAs Alguero:2023zol to calculate the relic density and DM-nucleon scattering cross-section after generating the model files using LanHEP Semenov:2014rea . The cyan-shaded region is ruled out by the most stringent direct search constraints from LZ-2022 LUX-ZEPLIN:2022qhg for heavier DM and CRESST-III CRESST:2019jnq and DarkSide-50 DarkSide:2018bpj for lighter DM. Clearly, we can see that the viable DM mass range that satisfies the correct relic density and direct detection rate ranges from around $500$ GeV to several TeV scales. Interestingly, future direct detection experiment Darwin DARWIN:2016hyl can probe a part of this high mass region up to $M_{\rm DM}\sim 3$ TeV.
In the discussions above, we have kept DM mass upto 10 TeV firstly due to the availability of direct detection prospects and secondly due to the fact that thermal DM gets overproduced if we go up by one order of magnitude $M_{\rm DM}\gtrsim\mathcal{O}(100)$ TeV, referred to as the unitarity bound Griest:1989wd ; Bhatia:2020itt . While such strict model-independent upper bound does not apply to non-thermal DM, we find such DM (fermion singlet) to be overproduced even with masses within tens of TeV, as mentioned earlier. Increasing fermion singlet DM masses beyond tens of TeV will further increase relic due to proportionate increase in Yukawa couplings from seesaw criteria and the fact that non-thermal DM yield increases with increasing interaction rate Hall:2009bx .
@@ V Observable $\Delta N_{\rm eff}$
at $2\sigma$ or $95\%$ C.L. including baryon acoustic oscillation (BAO) data. At $1\sigma$ C.L. it becomes more stringent to ${\rm N}_{\rm eff}=2.99\pm 0.17$. A similar bound also exists from big bang nucleosynthesis (BBN), $3.42.3<{\rm N}_{\rm eff}<$ 3.4 at $95\%$ C.L. Cyburt:2015mya . Recent observations by the DESI collaboration report $\Delta{\rm N}_{\rm eff}<0.4$ at $2\sigma$ DESI:2024mwx , which is weaker than the PLANCK bound given in Eq. (14). All these bounds are consistent with SM predictions ${\rm N^{SM}_{eff}}=3.045$ Mangano:2005cc ; Grohs:2015tfy ; deSalas:2016ztq . Future CMB experiment CMB-S4 is expected reach a much better sensitivity of $\Delta{\rm N}_{\rm eff}={\rm N}_{\rm eff}-{\rm N}^{\rm SM}_{\rm eff}=0.06$ Abazajian:2019eic , taking it closer to the SM prediction. Another future CMB experiment, specially, CMB-HD CMB-HD:2022bsz can probe $\Delta{\rm N}_{\rm eff}$ up to $0.014$ at $1\sigma$. Dirac nature of light neutrinos introduces additional relativistic degrees of freedom in the form of right chiral neutrinos, which can contribute to $\Delta{\rm N}_{\rm eff}$ either thermally or non-thermally. Light Dirac neutrino models often lead to enhanced $\Delta{\rm N}_{\rm eff}$, some recent works on which can be found in Abazajian:2019oqj ; FileviezPerez:2019cyn ; Nanda:2019nqy ; Han:2020oet ; Luo:2020sho ; Borah:2020boy ; Adshead:2020ekg ; Luo:2020fdt ; Mahanta:2021plx ; Du:2021idh ; Biswas:2021kio ; Borah:2022obi ; Li:2022yna ; Das:2023yhv . Depending upon the strength of Yukawa couplings associated with $\nu_{R}$, we can have either thermal or non-thermal contribution to $\Delta{\rm N}_{\rm eff}$ as we discuss in the following subsections.
V.1 Thermal $\Delta{\rm N}_{\rm eff}$
@@
We have three generations of right handed neutrinos ($\nu_{R_{i}}$) in our model, and their abundance in the Universe will be decided by the interaction strength and the masses of the particle with which they interact. The last three interaction terms of the Lagrangian (Eq. (II)) are the only three interactions that $\nu_{R}$ have with $N$ and $\chi$. The relative strength between the two types of Yukawa couplings $y_{i}$ and $y_{i}^{\prime}$ leads to three distinctive classes (Case-A, Case-B, Case-C) as we discussed before, in the context of neutrino mass and DM. With sufficiently large Yukawa coupling ($y_{i}^{\prime}$) as in Case-B, $\nu_{R}$ may enter the thermal bath with the SM degrees of freedom. As the Universe expands and the temperature falls, the scattering processes that kept $\nu_{R}$ in the thermal bath eventually go out of equilibrium, leading to the decoupling of $\nu_{R}$. For the combined scattering processes of $N_{L}\bar{\nu_{R}}\leftrightarrow N_{L}\bar{\nu_{R}}$ and $\chi\bar{\nu_{R}}\leftrightarrow\chi\bar{\nu_{R}}$, we check the decoupling or thermalization profile with two sets of benchmark parameters (masses and couplings) as shown in Fig. 6. As $N$ and $\chi$ masses lie at $\mathcal{O}$(TeV) range, it leads to the decoupling of $\nu_{R}$ from the thermal bath very early. Such early decoupling, above the electroweak scale, leads to a fixed value $\Delta{\rm N}_{\rm eff}=0.14$ for three generations of $\nu_{R}$. Clearly, this remains within reach of several future CMB experiments mentioned earlier.
V.2 Non-thermal $\Delta{\rm N}_{\rm eff}$
It is also possible to have observable $\Delta{\rm N}_{\rm eff}$ when $\nu_{R}$ does not enter the thermal bath. In such a case, $\nu_{R}$ can be produced non-thermally from the decay of the scalar $\chi$ as $\chi\rightarrow N\nu_{R_{i}}$ or from the decay of $N$ as $N\rightarrow\chi\nu_{R_{i}}$ depending upon our choice of fermion or scalar DM. Similar contributions can also arise from scatterings. With small Yukawa couplings for the $\bar{N_{L}}\chi\nu_{R}$ interaction (as in Case-A and Case-C), $\nu_{R}$ can not enter the thermal bath via the $N_{L}\bar{\nu_{R}}\leftrightarrow N_{L}\bar{\nu_{R}}$ (or $\chi\bar{\nu_{R}}\leftrightarrow\chi\bar{\nu_{R}}$) scattering process. In such a situation, where the scalar $\chi$ is in the thermal bath, $\nu_{R}$ are produced via the decay $\chi\rightarrow N\nu_{R}$ (assuming fermion DM scenario).
@@ V.2 Non-thermal $\Delta{\rm N}_{\rm eff}$
where $\Gamma_{\chi}$ denotes the partial decay width for the process $\chi\rightarrow N\nu_{R}$. The $\nu_{R}$ produced from $\chi$ decay will contribute to this additional radiation density. Therefore, with $\rho_{R}=\rho_{\rm SM}+\rho_{\nu_{R}}$, $\Delta{\rm N}_{\rm eff}$ for one species of $\nu_{R}$ (indicated by flavour index $\alpha$) calculated at the CMB epoch $(T_{\rm CMB})$ is
$\Delta{\rm N}^{\alpha}_{\rm eff}$ $=$ ${\rm N}^{\alpha}_{\rm eff}-{\rm N}_{\rm eff}^{\rm SM}$ (17)
@@
where, the extrapolation from the BBN epoch $(T_{\rm BBN})$ to CMB epoch $(T_{\rm CMB})$ is justified by the fact that both $\rho_{\nu_{L}},\rho_{\nu_{R}}$ redshift similarly for $T<T_{\rm BBN}$. One can simply calculate the energy density ratio at BBN epoch $\sim\mathcal{O}({10})$ MeV as $\nu_{R}$ production stops much earlier. The numerical factor 2 in Eq. (17) indicates the degrees of freedom of each $\nu_{R}$. The net contribution to $\Delta{\rm N}_{\rm eff}$ can be obtained by summing over flavour index $\alpha$ as $\Delta{\rm N}_{\rm eff}=\sum_{\alpha}\Delta{\rm N}^{\alpha}_{\rm eff}$. Ignoring the temperature dependence of the relativistic degrees of freedom, one can solve Eq. (15) analytically to find $\Delta{\rm N}_{\rm eff}$ Biswas:2022vkq .
@@
The decoupling profile for $\nu_{R}$ for two different benchmark choices of Yukawa coupling $y_{i}^{\prime}$ is shown in Fig. 6. For scalar ($\chi$) and fermion ($N$) masses in the TeV ballpark, we find that for $y_{i}^{\prime}\leq\mathcal{O}(10^{-3})$, the interaction $N_{L}\bar{\nu_{R}}\leftrightarrow N_{L}\bar{\nu_{R}}$ can not attain thermal equilibrium in the early Universe. Therefore, Case-A with $y_{i}^{\prime}\sim\mathcal{O}(10^{-11})$, $\nu_{R}$ will never enter the thermal bath, hence, thermal $\Delta{\rm N}_{\rm eff}$ is not possible. On the other hand, $\nu_{R}$ produced non-thermally from the decay of either $\chi$ or $N$ gives rise to negligible contribution to $\Delta{\rm N}_{\rm eff}$. As mentioned before, the Case-B, with $0.01\leq y_{1}^{\prime}\leq 0.1$, $0.03\leq y_{2}^{\prime}\leq 0.1$, $0.03\leq y_{3}^{\prime}\leq 0.1$, we have thermal contribution to $\Delta{\rm N}_{\rm eff}$. Due to the decoupling of $\nu_{R}$ much above the electroweak scale, we have a fixed contribution $\Delta{\rm N}_{\rm eff}=0.14$ for Case-B. Finally, in Case-C, Yukawa couplings are of the size $\mathcal{O}(10^{-6})$, and the corresponding interactions do not allow $\nu_{R}$ to enter the thermal bath. Considering the decay of $\chi\rightarrow\nu_{R}N$, we have shown the non-thermal contributions to $\Delta N_{\rm eff}$ in Fig. 7. Clearly, some part of the parameter space is already ruled out by PLANCK 2018 and DESI 2024 limits while most of the remaining parameter space remains within reach of future CMB experiments like SPT-3G, CMB-S4 and CMB-HD. As there are multiple $\mathcal{Z}_{2}$-odd scalars, it is possible to have $\chi_{i}\rightarrow\nu_{R}N$ type decays from one of the heavier scalars irrespective of fermion or scalar DM. While contribution to $\Delta{\rm N}_{\rm eff}$ from scalar decay can be calculated simply by considering the scalar $\chi_{i}$ to be in equilibrium, contributions from the decay $N\rightarrow\nu_{R}\chi_{j}$ will be more involved as the number density of non-thermal $N$ needs to be found by solving the respective Boltzmann equation Biswas:2022vkq .
@@
To summarise, observable $\Delta{\rm N}_{\rm eff}$ is possible in Case-B and Case-C. However, fermion DM is not possible in these two cases irrespective of the freeze-out or freeze-in mechanism. On the other hand, fermion DM produced via freeze-out mechanism is viable in Case-A even though the $\Delta{\rm N}_{\rm eff}$ remains out of reach from future CMB experiments. This conflict between observable $\Delta{\rm N}_{\rm eff}$ and DM relic can be avoided if we consider scalar DM. Scalar DM allows light scalars in the sub-TeV ballpark with observable direct detection rates while keeping collider prospects more promising compared to the fermion DM scenario where all $\mathcal{Z}_{2}$-odd particles remain in the multi-TeV ballpark.
It should be noted that we have calculated DM relic and $\Delta N_{\rm eff}$ for the points shown in Fig. 2, 3, 4. Since some of these points are disallowed by relic and $\Delta N_{\rm eff}$ constraints, shown in Fig. 5 and Fig. 7 respectively, it is possible, in principle, to rule out some of the points in Fig. 2, 3, 4 as well. However, as mentioned while describing the fit to neutrino data, we considered the entire loop function as a parameter instead of the details of scalar masses and mixing. On the other hand DM and $N_{\rm eff}$ phenomenology consider specific scalar DM masses satisfying those loop functions. Therefore, implementing the constraints from DM and $\Delta N_{\rm eff}$ on neutrino data allowed points in Fig. 2, 3, 4 will involve simultaneous generation of the complete scalar and fermion singlet mass spectra which we have not performed during our numerical scan. While a complete parameter scan using all phenomenological bounds is beyond the scope of the present work, we show the interplay of different bounds and observables highlighting the detection prospects of this Dirac scoto-seesaw model.